OpenStudy (bigbosssaint21):

An expression is shown below: 3x3y + 12xy - 3x2y - 12y Part A: Rewrite the expression so that the GCF is factored completely.

OpenStudy (neer2890):

3x3y+12xy-3x2y-12y =9xy+12xy-6xy-12y =21xy-6xy-12y =15xy-12y =3y(5x-4)

OpenStudy (bigbosssaint21):

Rewrite the expression completely factored.

OpenStudy (neer2890):

3y(5x-4)

OpenStudy (bigbosssaint21):

If the two middle terms were switched so that the expression became 3x^3y - 3x^2y + 12xy - 12y, would the factored expression no longer be equivalent to your answer in part B?

OpenStudy (neer2890):

ok.sorry i misunderstood the question.

OpenStudy (neer2890):

=3y(x^3+x^2+4x-4) =3y(x^2(x-1)+4(x-1)) =3y(x^2+4)(x-1)

OpenStudy (bigbosssaint21):

Write the expression x2 + 7x + 12 as a product of two linear expressions.

OpenStudy (neer2890):

=x^2+3x+4x+12 =x(x+3)+4(x+3) (x+3)(x+4)

OpenStudy (bigbosssaint21):

Write the expression x^2 + 7x + 12 as a product of two linear expressions.

OpenStudy (neer2890):

(x+3)(x+4) that's what i wrote

OpenStudy (bigbosssaint21):

lol sorry hang on

OpenStudy (bigbosssaint21):

Rewrite x^2 + 4x + 4 as a square of a linear expression

OpenStudy (bigbosssaint21):

@neer2890 , i got it typed write lol

OpenStudy (neer2890):

(x+2)(x+2)

OpenStudy (bigbosssaint21):

Do the expressions in parts A and B have a common factor? a is Write the expression x2 + 7x + 12 as a product of two linear expressions and b is Rewrite x2 + 4x + 4 as a square of a linear expression

OpenStudy (neer2890):

(x+2)^2

OpenStudy (neer2890):

no common factor.you can see

OpenStudy (bigbosssaint21):

The area of a square is (25x^2 - 20x + 4) square units. Determine the length of each side of the square by factoring the area expression completely

OpenStudy (neer2890):

you can find the roots of quadratic equation \[ax ^{2}+bx+c=0\] by \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (bigbosssaint21):

The area of a rectangle is (49x2 - 36y2) square units. Determine the dimensions of the rectangle by factoring the area expression completely.

OpenStudy (bigbosssaint21):

@neer2890 were almost done lol

OpenStudy (neer2890):

Area of rectangle= length* breadth (49x^2-36y^2) =(7x)^2-(6y)^2=(7x+6y)(7x-6y)

OpenStudy (bigbosssaint21):

The volume of a rectangular box is (x^3 + 2x^2 - 9x - 18) cubic units. Determine the dimensions of the rectangular box by factoring the volume expression completely

OpenStudy (neer2890):

(x+2) is a factor.

OpenStudy (neer2890):

so divide by (x+2)

OpenStudy (neer2890):

(x+2)(x^2-9) =(x+2)(x+3)(x-3)

OpenStudy (neer2890):

now i'm done buddy..lol

OpenStudy (bigbosssaint21):

for the night oooorrr that prob lol

OpenStudy (bigbosssaint21):

theres only 5 after this one @neer2890 f(x) = -16x2 + 60x + 16 Part A: What are the x-intercepts of the graph of the f(x)?

OpenStudy (neer2890):

i'm sorry, tired of doing the same problems..! don't mind please.

OpenStudy (bigbosssaint21):

your fine! thank you SOOOOOOO much though =)

OpenStudy (bigbosssaint21):

@patrickthestar can you help?

OpenStudy (bigbosssaint21):

f(x) = -16x^2 + 60x + 16 Part A: What are the x-intercepts of the graph of the f(x)?

OpenStudy (anonymous):

I have no idea man I suck at math

OpenStudy (bigbosssaint21):

lol

OpenStudy (bigbosssaint21):

@goformit100 ?

OpenStudy (bigbosssaint21):

f(x) = -16x^2 + 60x + 16 Part A: What are the x-intercepts of the graph of the f(x)?

OpenStudy (anonymous):

the answer is C purple because aliens don't wear hats

OpenStudy (bigbosssaint21):

for real though lol

OpenStudy (anonymous):

lol idk man im a sophomore and I still suck at math XD

OpenStudy (bigbosssaint21):

IM A SENIOR

OpenStudy (bigbosssaint21):

@wio ? help

OpenStudy (the_fizicx99):

This is algebra I, seniors are taking Calculus... For the x intercepts you can factor and set it equal to zero or use the Quadratic formula. BTW, these look awfully familiar to FLVS final exams. We aren't allowed to help you on exams.

OpenStudy (bigbosssaint21):

its certainly not a final, i wouldnt use openstudy for that, but yea i am in a crzy pickle where i was told a week from grad that i was missing agl 1 and they told me the year before that that i had everything i needed and just to chill, so ive been goin crazy

OpenStudy (the_fizicx99):

Hm, well doesn't matter. Factor it to find the x intercepts. Or use the Quad formula